OpenStudy (anonymous):

x/4-7>-10 will medal!!!please help

OpenStudy (anonymous):

You want to solve for the values of x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay you have \(\frac{x}{4}-7>-10\) the first step is the add 7 to both sides of the > sign, can you do that?

OpenStudy (anonymous):

so x/11>-3

OpenStudy (anonymous):

Hmm nope you have the right side correct, but when you add 7 to the left side it cancels out with the -7 to leave just \(\frac{x}{4}\)

OpenStudy (anonymous):

i tried to do it and i came up with x>-17

OpenStudy (anonymous):

\(\frac{x}{4}>-3\) now multiply both sides by 4 to get your answer

OpenStudy (anonymous):

oh so it would be x>-12

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

thanks